OpenStudy (anonymous):

The diagonals of a parallelogram have lengths 8 and 14 and they meet at a 60 degree angle. Find the area and the perimeter of the parallelogram.

OpenStudy (anonymous):

Use the Law of sines to find the area of one small triangle inside and multiply by 4

OpenStudy (anonymous):

\[ \frac{4}{2} 4 ( 7) \sin (60 {}^{\circ})=28 \sqrt{3} \]

OpenStudy (anonymous):

is that the area?

OpenStudy (anonymous):

yes